# Enumerating Finite Braid Group Orbits on $SL_{2}(\mathbb{C})$-Character Varieties

### v2

@@ 5.3.2 $G_{30}$
The group $G_{30}$ of order 14400 is also denoted $W(H_{4})$ with degrees $\deg(G_{30})=\{2,12,20,30\}$ [21, Table D.3]. All reflections in $G_{30}$ have nontrivial eigenvalue -1 and are conjugate to each other. Looking at the degrees, we extend the base field to $\mathbb{Q}(\zeta_{60})$. We found three distinct types of nice 4-tuples, the details are in Table 21.
@@
A $(-1,-1,-1,-1)$ $\{\frac{3}{10},\frac{3}{10},\frac{7}{10},\frac{7}{10}\}$
B $(-1,-1,-1,-1)$ $\left\{\frac{1}{6},\frac{1}{6},\frac{5}{6},\frac{5}{6}\right\}$
C $(-1,-1,-1,-1)$ $\{\frac{1}{10},\frac{1}{10},\frac{9}{10},\frac{9}{10}\}$
@@
For all 5 orbits, once we put the matrices in $SL_{2}(\mathbb{C})$, we see that the last matrix is the identity, hence these actually corresponds to orbits of 4 matrices multiplying to the identity. We compared these to the existing classification(s) by Lisovy-Tykhyy and Tykhyy ([22], [31]) as we did for the rank 3 complex reflection groups. Within each type, the two orbits are equivalent to the same orbit of $\mathcal{M}^{(4)}$ in Tykhyy’s classification, so we only provide the details once in Table 22.
@@
B $z^{10}$ $(2,2,2,-1)$ 31 $(\frac{2}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ $(0,\frac{1}{5},\frac{3}{5},\frac{3}{5})$ 96 18 7 $\langle 120,5\rangle$
@@ 5.3.3 $G_{32}$
The group $G_{32}$ of order 155520 is also denoted $W(L_{4})$ [21, Table D.3]. In this group, the nontrivial eigenvalues of reflections are in $\{\zeta_{3},\zeta_{3}^{2}\}$ corresponding to two conjugacy classes. For $G_{32}$, $\deg(G_{32})=\{12,18,24,30\}$ so the eigenvalues of multiplicity 2 must be 12-th roots of unity. Expanding the field to $\mathbb{Q}(\zeta_{12})$ we found 5 distinct eigenvalue sets for the inverse product. Note that to ensure we found all examples, we would have to run the code over $\mathbb{Q}(\zeta_{360})$ since there may be matrices with one or two eigenvalues other than the 12-th roots which would not be found, however we claim it is enough to check over $\mathbb{Q}(\zeta_{12})$.
@@
If $A=A_{1}A_{2}A_{3}A_{4}$, with $[A_{1},A_{2},A_{3},A_{4}]$ a nice tuple in $G_{32}$ so that $A$ has a nontrivial eigenvalue of multiplicity 2, then the other two eigenvalues of $A$ are both $m$-th roots of unity (not necessarily primitive) for the same $m\in\{12,18,24,30\}$.
@@ C $(w^{2},w^{2},w^{2},w^{2})$$\{\frac{1}{12},\frac{1}{12},\frac{7}{12},\frac{7}{12}\}$
The middle convolution details can be found in Table 36. Different exemplars nor the inverse types produce new orbits. The two orbits from Type B are distinct as are the two orbits from Type C. Of the five middle convolutions computed in $G_{32}$, after putting the matrices into $SL_{2}(\mathbb{C})$ three of them (one from Type B and both orbits from Type C) produce tuples where the 5-th matrix is the identity. For these three, we compare with the classification(s) by Lisovy-Tykhyy and Tykhyy ([22], [31]) to see which $\mathcal{M}^{(4)}$ orbits they correspond. The orbits from Type C correspond to the same orbit in Tykhyy’s classification so we only provide the details for it once. The tuples that correspond to nontrivial local systems on the 5-punctured sphere are those coming from Type A and the other middle convolution from Type B. For the Type B tuple, after putting the matrices into $SL_{2}(\mathbb{C})$, we get that the product of the third and fourth matrix is not of finite order, hence we cannot compute its signature to compare with Tykhyy’s table of $\mathcal{M}^{(5)}$. The matrices in this tuple $MC_{-z^{2}+1}(\mathbf{B})$ are:
@@
B $z^{2}$ $(0,3,0,-2)$ Type II, $X^{\prime}=1,X^{\prime\prime}=2$ $(\frac{5}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},1,\frac{2}{3})$ 3 2 60 $\langle 24,4\rangle$
C $z$ $(4,4,4,-8)$ Type IV, $\omega=4$ $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 6 4 20 $\langle 48,28\rangle$
@@
A $-z^{2}+1$ $(\frac{1}{2},\frac{5}{6},\frac{1}{6},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},1,\frac{1}{3},\frac{1}{2},\frac{5}{6},\frac{1}{2},1)$ 1 4 40 $\langle 24,4\rangle$
B $-z^{2}+1$ $\geq 350$ 0
@@ 5.4.1 $G_{32}$
A $(w,w,-w-1,w,w)$$\{\frac{1}{2},\frac{5}{6},\frac{5}{6},\frac{5}{6}\}$
B $(-w-1,w,-w-1,w,w)$$\left\{\frac{1}{6},\frac{5}{6},\frac{5}{6},\frac{5}{6}\right\}$
@@
The details of the middle convolution for both types is presented in Table 37. The $SL_{2}(\mathbb{C})$ subgroups generated by both tuple types is the dicyclic group of order 24. Inspecting the 6 $SL_{2}(\mathbb{C})$ matrices we get for both of these types, both tuples satisfy the description of ‘TYPE’ A of $\mathcal{M}^{(6)}$ belonging to the dihedral group in Tykhyy’s classification [31]. In these cases we also could not check if different examplars/ inverses produce different orbits as we could not generate the list of all nice tuples.
@@ Appendix A Middle Convolution Tables For Nice 3-Tuples
$\frac{1}{2}$ $-1$ $\begin{pmatrix}1&0\\ -q^{3}-q^{2}+1&1\end{pmatrix}$ $\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$ $(1,1,1,1)$ 40 0
$\frac{1}{10}$ $-q^{3}$ $\begin{pmatrix}q^{3}&0\\ -q^{3}&1\end{pmatrix}$ $\begin{pmatrix}q^{3}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ q^{3}&q^{3}\end{pmatrix}$ $(q,q,q,q^{2})$ 10 600
$\frac{9}{10}$ $-q^{2}$ $\begin{pmatrix}q^{2}&0\\ -q^{2}&1\end{pmatrix}$ $\begin{pmatrix}q^{2}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ q^{2}&q^{2}\end{pmatrix}$ $(q^{4},q^{4},q^{4},q^{3})$ 10 600
@@
$\frac{1}{2}$ $-1$ $\small{\begin{pmatrix}1&0\\ 1&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ q^{3}+q^{2}+2&1\end{pmatrix}}$ $(1,1,1,1)$ 40 0
$\frac{3}{10}$ $-q^{4}$ $\small{\begin{pmatrix}q^{4}&0\\ -q^{4}+q^{3}+1&1\end{pmatrix}}$ $\small{\begin{pmatrix}q^{4}&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ q^{4}-q^{3}-1&q^{4}\end{pmatrix}}$ $(q^{3},q^{3},q^{3},q)$ 10 600
$\frac{7}{10}$ $-q$ $\small{\begin{pmatrix}q&0\\ q^{2}-q+1&1\end{pmatrix}}$ $\small{\begin{pmatrix}q&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -q^{2}+q-1&q\end{pmatrix}}$ $(q^{2},q^{2},q^{2},q^{4})$ 10 600
@@
$\frac{1}{2}$ $-1$ $\small{\begin{pmatrix}u-1&3u-5\\ 1&-u+3\end{pmatrix}}$ $\small{\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ u+1&1\end{pmatrix}}$ $(1,1,1,1)$ 72 0
@@
$\frac{5}{6}$ $-z^{10}$ $\small{\begin{pmatrix}z^{5}+v&z^{10}\\ -u-2v-z^{5}&v\end{pmatrix}}$ $\small{\begin{pmatrix}z^{10}&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ v+z^{10}&z^{10}\end{pmatrix}}$ $\small{(z^{10},z^{10},z^{10},1)}$ 18 360
[…]
tableHere $q=\zeta_{5},z=\zeta_{30}$, $u=z^{7}-z^{3}-z^{2}+1$ and $v=-z^{4}+z$
@@
$\frac{1}{14}$ $-z^{4}$ $\small{\begin{pmatrix}z^{5}+z^{4}+z+1&-z^{5}+z^{4}+z^{3}+1\\ \>-z^{2}&-z^{5}-z\end{pmatrix}}$ $\small{\begin{pmatrix}z^{4}&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{4}&z^{4}\end{pmatrix}}$ $(z^{5},z^{5},z^{5},z^{6})$ 28 0
$\frac{9}{14}$ $-z$ $\small{\begin{pmatrix}-z^{6}-z^{5}-z^{4}&z^{6}-z^{3}+1\\ -z^{4}&-z^{3}-z^{2}\end{pmatrix}}$ $\small{\begin{pmatrix}z&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z&z\end{pmatrix}}$ $(z^{3},z^{3},z^{3},z^{5})$ 28 0
$\frac{11}{14}$ $-z^{2}$ $\small{\begin{pmatrix}-z^{5}-z^{3}-z&-z^{6}+z^{5}+z^{2}+1\\ -z&-z^{6}-z^{4}\end{pmatrix}}$ $\small{\begin{pmatrix}z^{2}&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{2}&z^{2}\end{pmatrix}}$ $-(z^{6},z^{6},z^{6},z^{3})$ 28 0
[…]
tableHere $z=\zeta_{7}$
@@
$\frac{5}{6}$ $-w$ $\begin{pmatrix}1&-w\\ 0&-w^{2}\end{pmatrix}$ $\begin{pmatrix}-w^{2}&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-1\end{pmatrix}$ $(iw^{2},iw^{2},i,iw^{2})$ 12 72
$\frac{1}{6}$ $-w^{2}$ $\begin{pmatrix}1&-w\\ 0&-1\end{pmatrix}$ $\begin{pmatrix}-1&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -w&-w\end{pmatrix}$ $(i,i,iw,iw^{2})$ 12 72
@@
$\frac{7}{9}$ $-z^{10}$ $\small{\begin{pmatrix}z^{2}&-z^{8}+z^{6}\\ z^{16}-z^{14}&z^{4}-z^{2}+1\end{pmatrix}}$ $\small{\begin{pmatrix}u&z^{6}\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{10}&u\end{pmatrix}}$ $\small{(u,-z^{10},-z^{10},1)}$ 36 0
@@
$\frac{11}{12}$ $-z^{9}+z^{3}$ $\frac{1}{2}\cdot\begin{pmatrix}z^{9}+1&-z^{9}+1\\ -z^{9}+1&z^{9}+1\end{pmatrix}$ $\begin{pmatrix}z^{9}&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{9}\end{pmatrix}$ $(q^{63},q^{63},q^{63},q^{27})$ 4 96
@@
$\frac{2}{3}$ $-z^{6}$ $\begin{pmatrix}1&z^{6}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-1&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-1&1\end{pmatrix}$ $(1,1,1,1)$ 16 0
@@
$\frac{7}{12}$ $-z^{3}$ $\begin{pmatrix}z^{9}&-z^{9}+1\\ u&-u+1\end{pmatrix}$ $\begin{pmatrix}z^{9}&1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{9}-z^{3}&z^{3}\end{pmatrix}$ $(q^{3},q^{27},q^{33},q^{9})$ 24 0
$\frac{5}{6}$ $-z^{6}+1$ $\begin{pmatrix}1&0\\ -z^{6}+1&z^{6}\end{pmatrix}$ $\begin{pmatrix}-1&1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{6}+2&z^{6}-1\end{pmatrix}$ $(z^{15},z^{9},z^{6}-1,1)$ 24 72
@@
$\frac{11}{18}$ $-z^{4}$ $\begin{pmatrix}-z^{16}&0\\ z^{16}-z^{14}+1&1\end{pmatrix}$ $\begin{pmatrix}-z^{16}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{16}+z^{4}&z^{4}\end{pmatrix}$ $(z,z,z^{16},-1)$ 36 0
$\frac{17}{18}$ $-z^{16}$ $\begin{pmatrix}z^{10}&0\\ -z^{10}-z^{2}+1&1\end{pmatrix}$ $\begin{pmatrix}z^{10}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{16}+z^{10}&z^{16}\end{pmatrix}$ $(z^{13},z^{13},z^{10},1)$ 36 0
@@
$\frac{5}{6}$ $-z^{6}+1$ $\begin{pmatrix}z^{6}&z^{6}\\ -2z^{6}+1&-z^{6}\end{pmatrix}$ $\begin{pmatrix}-1&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-2&z^{6}-1\end{pmatrix}$ $(i,i,w,w+1)$ 12 18
$\frac{1}{6}$ $z^{6}$ $(-z^{6}+1)$ $(-z^{6}+1)$ $(-z^{6})$

### v3

@@ 5.3.2 $G_{30}$
The group $G_{30}$ of order 14400 is also denoted $W(H_{4})$ with degrees $\deg(G_{30})=\{2,12,20,30\}$ [ld2009, Table D.3]. All reflections in $G_{30}$ have nontrivial eigenvalue -1 and are conjugate to each other. Looking at the degrees, we extend the base field to $\mathbb{Q}(\zeta_{60})$. We found three distinct types of nice 4-tuples, the details are in Table 25.
@@
A (1,1,1,1)(-1,-1,-1,-1) $\{\frac{3}{10},\frac{3}{10},\frac{7}{10},\frac{7}{10}\}$
B (1,1,1,1)(-1,-1,-1,-1) $\left\{\frac{1}{6},\frac{1}{6},\frac{5}{6},\frac{5}{6}\right\}$
C (1,1,1,1)(-1,-1,-1,-1) $\{\frac{1}{10},\frac{1}{10},\frac{9}{10},\frac{9}{10}\}$
@@
For all 5 orbits, once we put the matrices in $SL_{2}(\mathbb{C})$, we see that the last matrix is the identity, hence these actually corresponds to orbits of 4 matrices multiplying to the identity. We compared these to the existing classification(s) by Lisovyy-Tykhyy and Tykhyy ([T-L], [Tyk]) as we did for the rank 3 complex reflection groups. Within each type, the two orbits are equivalent to the same orbit of $\mathcal{M}^{(4)}$ in Tykhyy’s classification, so we only provide the details once in Table 26.
@@
B $z^{10}$ (2,2,2,1)(2,2,2,-1) 31 $(\frac{2}{3},\frac{2}{3},\frac{2}{3},\frac{2}{3})$ $(0,\frac{1}{5},\frac{3}{5},\frac{3}{5})$ 96 18 7 $\langle 120,5\rangle$
@@ 5.3.3 $G_{32}$
The group $G_{32}$ of order 155520 is also denoted $W(L_{4})$ [ld2009, Table D.3]. In this group, the nontrivial eigenvalues of reflections are in $\{\zeta_{3},\zeta_{3}^{2}\}$ corresponding to two conjugacy classes. For $G_{32}$, $\deg(G_{32})=\{12,18,24,30\}$ so the eigenvalues of multiplicity 2 must be 12-th roots of unity. Expanding the field to $\mathbb{Q}(\zeta_{12})$ we found 5 distinct eigenvalue sets for the inverse product. Note that to ensure we found all examples, we would have to run the code over $\mathbb{Q}(\zeta_{360})$ since there may be matrices with one or two eigenvalues other than the 12-th roots which would not be found, however we claim it is enough to check over $\mathbb{Q}(\zeta_{12})$.
@@
If $A=A_{1}A_{2}A_{3}A_{4}$, with $A_{1},A_{2},A_{3},A_{4}]$ a nice tuple in $G_{32}$ so that $A$ has a nontrivial eigenvalue of multiplicity 2, then the other two eigenvalues of $A$ are both $m$-th roots of unity (not necessarily primitive) for the same $m\in\{12,18,24,30\}$.
@@ C $(w^{2},w^{2},w^{2},w^{2})$ $\{\frac{1}{12},\frac{1}{12},\frac{7}{12},\frac{7}{12}\}$
The middle convolution details can be found in Table 44. Different exemplars nor the inverse types produce new orbits. The two orbits from Type B are distinct as are the two orbits from Type C. Of the five middle convolutions computed in $G_{32}$, after putting the matrices into $SL_{2}(\mathbb{C})$ three of them (one from Type B and both orbits from Type C) produce tuples where the 5-th matrix is the identity. For these three, we compare with the classification(s) by Lisovyy-Tykhyy and Tykhyy ([T-L], [Tyk]) to see which $\mathcal{M}^{(4)}$ orbits they correspond. The orbits from Type C correspond to the same orbit in Tykhyy’s classification so we only provide the details for it once. The tuples that correspond to nontrivial local systems on the 5-punctured sphere are those coming from Type A and the other middle convolution from Type B. For the Type B tuple, after putting the matrices into $SL_{2}(\mathbb{C})$, we get that the product of the third and fourth matrix is not of finite order, hence we cannot compute its signature to compare with Tykhyy’s table of $\mathcal{M}^{(5)}$. The matrices in this tuple $MC_{-z^{2}+1}(\mathbf{B})$ are:
@@
B $z^{2}$ (0,3,0,2)(0,3,0,-2) Type II, $X^{\prime}=1,X^{\prime\prime}=2$ $(\frac{5}{6},\frac{1}{2},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},\frac{1}{2},1,\frac{2}{3})$ 3 2 60 $\langle 24,4\rangle$
C $z$ (4,4,4,8)(4,4,4,-8) Type IV, $\omega=4$ $(\frac{1}{4},\frac{1}{4},\frac{1}{4},\frac{1}{4})$ $(\frac{1}{3},\frac{1}{3},\frac{1}{3},0)$ 6 4 20 $\langle 48,28\rangle$
@@
A $z^{2}+1$ $(\frac{1}{2},\frac{5}{6},\frac{1}{6},\frac{1}{6},\frac{1}{2})$ $(\frac{1}{2},1,\frac{1}{3},\frac{1}{2},\frac{5}{6},\frac{1}{2},1)$ 1 4 40 $\langle 24,4\rangle$
B $z^{2}+1 -$ - - - $\geq 350$ 0
@@ 5.4.1 $G_{32}$
A (w,w,w1,w,w)($w,$-w-1,w,w) $\{\frac{1}{2},\frac{5}{6},\frac{5}{6},\frac{5}{6}\}$
B (w1,w,w1,w,w)(-w-1,w,-w-1,w,w) $\left\{\frac{1}{6},\frac{5}{6},\frac{5}{6},\frac{5}{6}\right\}$
@@
The details of the middle convolution for both types is presented in Table 45. The $SL_{2}(\mathbb{C})$ subgroups generated by both tuple types is the dicyclic group of order 24. Inspecting the 6 $SL_{2}(\mathbb{C})$ matrices we get for both of these types, both tuples satisfy the description of ‘TYPE’ A of $\mathcal{M}^{(6)}$ belonging to the dihedral group in Tykhyy’s classification [Tyk]. In these cases we also could not check if different examplars/ inverses produce different orbits as we could not generate the list of all nice tuples.
[…]
The notation for all the appendices are the same. The $\xi$’ column gives the residue of the root of unity in the $\lambda$’ column used to compute middle convolution. The ‘Character’ column is the character we tensor our tuple with to get matrices in $SL_{2}(\mathbb{C})$. The column ‘O. Size’ is the size of the $SL_{2}(\mathbb{C})$ character variety orbit of the tuple (i.e after it is tensored with the character). In these tables ‘S. Size’ denotes the size of the $GL_{2}(\mathbb{C})$ subgroup generated by the $2\times 2$ matrices we get after middle convolution but before we tensor with the character. The $SL_{2}(\mathbb{C})$ subgroup size after tensoring with the character is provided in each section for each group as part of the comparison tables. For infinite subgroups, the order is listed as ‘0’, following the notation used in Magma.
@@ Appendix A Middle Convolution Tables For Nice 3-Tuples
$\frac{1}{2}$ 1-1 $\begin{pmatrix}1&0\\ -q^{3}-q^{2}+1&1\end{pmatrix}$ $\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 1&1\end{pmatrix}$ $(1,1,1,1)$ 40 0
$\frac{1}{10}$ $q^{3}$ $\begin{pmatrix}q^{3}&0\\ -q^{3}&1\end{pmatrix}$ $\begin{pmatrix}q^{3}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ q^{3}&q^{3}\end{pmatrix}$ $(q,q,q,q^{2})$ 10 600
$\frac{9}{10}$ $q^{2}$ $\begin{pmatrix}q^{2}&0\\ -q^{2}&1\end{pmatrix}$ $\begin{pmatrix}q^{2}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ q^{2}&q^{2}\end{pmatrix}$ $(q^{4},q^{4},q^{4},q^{3})$ 10 600
@@
$\frac{1}{2}$ 1-1 $\small{\begin{pmatrix}1&0\\ 1&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ q^{3}+q^{2}+2&1\end{pmatrix}}$ $(1,1,1,1)$ 40 0
$\frac{3}{10}$ $q^{4}$ $\small{\begin{pmatrix}q^{4}&0\\ -q^{4}+q^{3}+1&1\end{pmatrix}}$ $\small{\begin{pmatrix}q^{4}&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ q^{4}-q^{3}-1&q^{4}\end{pmatrix}}$ $(q^{3},q^{3},q^{3},q)$ 10 600
$\frac{7}{10}$ q-q $\small{\begin{pmatrix}q&0\\ q^{2}-q+1&1\end{pmatrix}}$ $\small{\begin{pmatrix}q&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -q^{2}+q-1&q\end{pmatrix}}$ $(q^{2},q^{2},q^{2},q^{4})$ 10 600
@@
$\frac{1}{2}$ 1-1 $\small{\begin{pmatrix}u-1&3u-5\\ 1&-u+3\end{pmatrix}}$ $\small{\begin{pmatrix}1&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ u+1&1\end{pmatrix}}$ $(1,1,1,1)$ 72 0
@@
$\frac{5}{6}$ $z^{10}$ $\small{\begin{pmatrix}z^{5}+v&z^{10}\\ -u-2v-z^{5}&v\end{pmatrix}}$ $\small{\begin{pmatrix}z^{10}&-1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ v+z^{10}&z^{10}\end{pmatrix}}$ $\small{(z^{10},z^{10},z^{10},1)}$ 18 360
Table 31: Here $q=\zeta_{5},z=\zeta_{30}$, $u=z^{7}-z^{3}-z^{2}+1$ and $v=-z^{4}+z$
@@
$\frac{1}{14}$ $z^{4}$ $\small{\begin{pmatrix}z^{5}+z^{4}+z+1&-z^{5}+z^{4}+z^{3}+1\\ \>-z^{2}&-z^{5}-z\end{pmatrix}}$ $\small{\begin{pmatrix}z^{4}&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{4}&z^{4}\end{pmatrix}}$ $(z^{5},z^{5},z^{5},z^{6})$ 28 0
$\frac{9}{14}$ z-z $\small{\begin{pmatrix}-z^{6}-z^{5}-z^{4}&z^{6}-z^{3}+1\\ -z^{4}&-z^{3}-z^{2}\end{pmatrix}}$ $\small{\begin{pmatrix}z&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z&z\end{pmatrix}}$ $(z^{3},z^{3},z^{3},z^{5})$ 28 0
$\frac{11}{14}$ $z^{2}$ $\small{\begin{pmatrix}-z^{5}-z^{3}-z&-z^{6}+z^{5}+z^{2}+1\\ -z&-z^{6}-z^{4}\end{pmatrix}}$ $\small{\begin{pmatrix}z^{2}&1\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{2}&z^{2}\end{pmatrix}}$ $(z^{6},z^{6},z^{6},z^{3})$ 28 0
Table 32: Here $z=\zeta_{7}$
@@
$\frac{5}{6}$ w-w $\begin{pmatrix}1&-w\\ 0&-w^{2}\end{pmatrix}$ $\begin{pmatrix}-w^{2}&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -1&-1\end{pmatrix}$ $(iw^{2},iw^{2},i,iw^{2})$ 12 72
$\frac{1}{6}$ $w^{2}$ $\begin{pmatrix}1&-w\\ 0&-1\end{pmatrix}$ $\begin{pmatrix}-1&w\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -w&-w\end{pmatrix}$ $(i,i,iw,iw^{2})$ 12 72
@@
$\frac{7}{9}$ $z^{10}$ $\small{\begin{pmatrix}z^{2}&-z^{8}+z^{6}\\ z^{16}-z^{14}&z^{4}-z^{2}+1\end{pmatrix}}$ $\small{\begin{pmatrix}u&z^{6}\\ 0&1\end{pmatrix}}$ $\small{\begin{pmatrix}1&0\\ -z^{10}&u\end{pmatrix}}$ $\small{(u,-z^{10},-z^{10},1)}$ 36 0
@@
$\frac{11}{12}$ $z^{9}+z^{3}$ $\frac{1}{2}\cdot\begin{pmatrix}z^{9}+1&-z^{9}+1\\ -z^{9}+1&z^{9}+1\end{pmatrix}$ $\begin{pmatrix}z^{9}&0\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ 0&z^{9}\end{pmatrix}$ $(q^{63},q^{63},q^{63},q^{27})$ 4 96
@@
$\frac{2}{3}$ $z^{6}$ $\begin{pmatrix}1&z^{6}\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-1&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-1&1\end{pmatrix}$ $(1,1,1,1)$ 16 0
@@
$\frac{7}{12}$ $z^{3}$ $\begin{pmatrix}z^{9}&-z^{9}+1\\ u&-u+1\end{pmatrix}$ $\begin{pmatrix}z^{9}&1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{9}-z^{3}&z^{3}\end{pmatrix}$ $(q^{3},q^{27},q^{33},q^{9})$ 24 0
$\frac{5}{6}$ $z^{6}+1$ $\begin{pmatrix}1&0\\ -z^{6}+1&z^{6}\end{pmatrix}$ $\begin{pmatrix}-1&1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{6}+2&z^{6}-1\end{pmatrix}$ $(z^{15},z^{9},z^{6}-1,1)$ 24 72
@@
$\frac{11}{18}$ $z^{4}$ $\begin{pmatrix}-z^{16}&0\\ z^{16}-z^{14}+1&1\end{pmatrix}$ $\begin{pmatrix}-z^{16}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ -z^{16}+z^{4}&z^{4}\end{pmatrix}$ $(z,z,z^{16},-1)$ 36 0
$\frac{17}{18}$ $z^{16}$ $\begin{pmatrix}z^{10}&0\\ -z^{10}-z^{2}+1&1\end{pmatrix}$ $\begin{pmatrix}z^{10}&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{16}+z^{10}&z^{16}\end{pmatrix}$ $(z^{13},z^{13},z^{10},1)$ 36 0
@@
$\frac{5}{6}$ $z^{6}+1$ $\begin{pmatrix}z^{6}&z^{6}\\ -2z^{6}+1&-z^{6}\end{pmatrix}$ $\begin{pmatrix}-1&-1\\ 0&1\end{pmatrix}$ $\begin{pmatrix}1&0\\ z^{6}-2&z^{6}-1\end{pmatrix}$ $(i,i,w,w+1)$ 12 18
$\frac{1}{6}$ $z^{6}$ $(-z^{6}+1)$ $(-z^{6}+1)$ $(-z^{6})$ - - -